Anybody know how to solve this one
Given f(x)=-5x^4+x^3+2x^2+3x+1, and k=1, express f(x) in the form f(x)=(x-k)q(x)+r

thank you very much for your time

Answer by stanbon(75887)

Given f(x)=-5x^4+x^3+2x^2+3x+1, and k=1, express f(x) in the form f(x)=(x-k)q(x)+r
-----------
Using synthetic division you get:
f(x)=(x-1)(-5x^3-4x^2-2x+1)+2
Cheers,
Stan H.